# Nesterov’s Accelerated Jacobi-Type Methods for Large-scale Symmetric Positive Semidefinite Linear Systems

### v1

@@ 4.4 Scalability in parallel computing environments
We test the scalability of the distributed algorithm on the following six sparse matrices: G3_circuit, bone010, apache2, audikw_1, StocF-1465, and thermal2, which are SPD matrices from the SuiteSparse Matrix Collection. Various properties of these matrices used in the test are presented in Table 1. Load imbalance is defined as the ratio of the maximum number of nonzeros assigned to a processor to the average number of nonzeros in each processor: $\dfrac{p*\max_{i}nnz(Q^{i,:})}{nnz(Q)}$.
@@
Figure 6 summarizes the scalability test results where the distributed algorithm is run for 100 iterations. The cost of the construction of the matrix $J$ (”setup J”) is constant regardless of the number of iterations, and hence the construction cost is minor when the number of iterations is large. Solving local linear equations defined by $J^{i,i}$ (”solve J”) and matrix-vector multiplication (”matvec”) are scalable. Consequently, the entire distributed algorithm (”totals”) is scalable.
@@
Figure 6: Scalability of the distributed version of Algorithm 3 against up to 121 processors. The distributed algorithm is run for $100$ iterations.
@@ 5 Conclusions
In this paper, we have considered solving large-scale symmetric positive semidefinite (SPSD) systems of linear equations via Jacobi-type iterative methods, which improve the classical Jacobi-type iterative methods by applying the Nesterov’s acceleration technique. The proposed method has been shown to be convergent for any consistent SPSD systems with an $O\left(\frac{1}{t^{2}}\right)$ convergence rate in terms of objective function values associated with the underlying linear system, provided that the proximal term in the Nesterov’s framework is induced by a symmetric positive semidefinite matrix. We have also proposed and analyzed a restarted version of the former method and shown that it has an $O\left(\frac{(\log_{2}(t))^{2}}{t^{2}}\right)$ worst-case convergence rate, under the condition that the matrix $Q$ is positive definite. To evaluate the practical performance of the proposed methods, we have conducted extensive numerical experiments. Our numerical results indicate that the proposed method outperforms the classical Jacobi-type methods and the conjugate gradient method, and shares a comparable performance as the diagonally preconditioned conjugate gradient method. The speed-up tests also provide strong evidence demonstrating that the proposed method is highly scalable.
@@ Acknowledgments
Ling Liang and Haizhao Yang were partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, and the Office of Naval Research Award N00014-23-1-2007.

### v2

@@ 4.4 Scalability in parallel computing environments
We test the scalability of the distributed algorithm with a diagonal preconditioner on the following six sparse matrices: G3_circuit, bone010, apache2, audikw_1, StocF-1465, and thermal2, which are SPD matrices from the SuiteSparse Matrix Collection. Various properties of these matrices used in the test are presented in Table 2. Load imbalance is defined as the ratio of the maximum number of nonzeros assigned to a processor to the average number of nonzeros in each processor: $\dfrac{p\cdot\max_{i}\{{\rm nnz}(Q^{i,:})\}}{{\rm nnz}(Q)}$.
@@
Figure 5 summarizes the scalability test results where the distributed algorithm is run for 100 iterations. The cost of the construction of the matrix $J$ is constant regardless of the number of iterations, and the construction cost is minor when the number of iterations is large. Solving local linear equations defined by $J_{i}$ (“solve J”) and matrix-vector multiplication (“matvec”) are scalable. The restarting technique requires a dot product (“dot”) which is known to be not scalable. Fortunately, the dot product only needs to be computed periodically but not at each every iteration. Consequently, the entire distributed algorithm (“totals”) is scalable until the dot product becomes a scalability bottleneck. Note that this does not indicate the proposed algorithm is not competitive as dot products widely present in competitors including PCG and CG. In fact, Figure 6 shows that the proposed distributed algorithm (“p acc-jacobi”) admits excellent performance in terms of scalability, which is comparable to the parallel PCG (“KSPCG”).
@@
Figure 5: Scalability of the distributed version of Algorithm 3 against up to 32 processors. The distributed algorithm is run for $100$ iterations.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Comparisons between the proposed distributed algorithm and pcg with the BJACOBI preconditioner. Both algorithms run for $100$ iterations.
@@ 5 Conclusions
In this paper, we have considered solving large-scale symmetric positive semidefinite (SPSD) systems of linear equations via Jacobi-type iterative methods, which improve the classical Jacobi-type iterative methods by applying Nesterov’s acceleration technique. The proposed method has been shown to be convergent for any consistent SPSD systems with an $O\left(\frac{1}{t^{2}}\right)$ convergence rate in terms of objective function values associated with the underlying linear system, provided that the proximal term in the Nesterov’s framework is induced by a symmetric positive semidefinite matrix. We have also proposed and analyzed a restarted version of the former method and shown that it has an $O\left(\frac{(\log_{2}(t))^{2}}{t^{2}}\right)$ worst-case convergence rate, under the condition that the matrix $Q$ is positive definite. To evaluate the practical performance of the proposed methods, we have conducted extensive numerical experiments. Our numerical results indicate that the proposed method outperforms the classical Jacobi-type methods and the conjugate gradient method, and shares a comparable performance as the diagonally preconditioned conjugate gradient method. The speed-up tests also provide strong evidence demonstrating that the proposed method is highly scalable.
@@ Acknowledgments
Haizhao Yang was partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.
We thank the editor and three anonymous reviewers for providing valuable suggestions, which have helped to improve the quality of the paper.